# Foundations of Quantum Granular Computing with Effect-Based Granules, Algebraic Properties and Reference Architectures

### v1

@@ Abstract
This paper develops the foundations of Quantum Granular Computing (QGC), extending classical granular computing—including fuzzy, rough, and shadowed granules—to the quantum regime. Quantum granules are modeled as effects on a finite-dimensional Hilbert space, so that granular memberships are given by Born probabilities. This operator-theoretic viewpoint provides a common language for sharp (projective) and soft (non-projective) granules and embeds granulation directly into the standard formalism of quantum information theory. On this basis, we establish foundational results for effect-based quantum granules, including normalization and monotonicity properties, the emergence of Boolean islands from commuting families, granular refinement under Lüders updates, and the evolution of granules under quantum channels via the adjoint channel in the Heisenberg picture. We further connect QGC with quantum detection and estimation theory by interpreting the effect operators realizing Helstrom’s minimum-error measurement for binary state discrimination as Helstrom-type decision granules, i.e., soft quantum counterparts of Bayes-optimal decision regions. Building on these results, we introduce Quantum Granular Decision Systems (QGDS) together with three reference architectures: Measurement-Driven Granular Partitioning (MDGP), Variational Effect Learning (VEL), and Hybrid Classical–Quantum pipelines (HCQ). These architectures specify how quantum granules can be defined, learned, and integrated with classical components while remaining compatible with near-term quantum hardware. Case studies on qubit granulation, two-qubit parity effects, and Helstrom-style soft decisions illustrate how QGC reproduces fuzzy-like graded memberships and smooth decision boundaries while exploiting non-commutativity, contextuality, and entanglement. Overall, the framework provides a unified and mathematically grounded basis for operator-valued granules in quantum information processing, granular reasoning, and intelligent systems.
@@ 1 Introduction
Classical granular computing (GrC) provides a principled way to structure information through granules—coherent units such as fuzzy sets, rough sets, interval-valued descriptions, and shadowed regions—that enable reasoning under uncertainty and abstraction [41, 26]. These models play a central role in intelligent systems and have been expanded through multilevel approximation schemes and structured granulation methodologies [36, 17]. Further expressive power has been achieved through shadowed sets, interval type-2 fuzzy sets [27], and Mediative Fuzzy Logic (MFL) [24], which allow granules to encode partial compatibility, conflict, and higher-order uncertainty. Nonetheless, classical granules typically operate within Boolean or fuzzy lattices, where global compatibility is implicitly assumed, limiting their ability to represent phenomena driven by fundamentally incompatible observations.
@@
Quantum systems provide a natural setting in which these limitations become explicit. In quantum mechanics, states, measurements, and transformations are operators on a Hilbert space, and observable quantities generally do not commute [25, 10]. This induces contextuality: not all observables admit a joint description, and measurement can irreversibly disturb the system. At the same time, quantum information theory has shown that operator-valued objects such as effects, POVMs, and quantum channels offer a flexible language for soft decisions, uncertainty, and partial information [4, 10]. The lattice of closed subspaces is orthomodular and non-distributive [3], and the set of effects forms an effect algebra [9], indicating that quantum theory already contains native algebraic structures suitable for granular modeling.
@@
We, therefore, adopt the viewpoint that quantum granules should be modeled as effects, i.e., positive operators bounded by the identity. Effects are standard generalized events in quantum information, providing soft acceptance regions parameterized by their spectrum. Their associated Born probabilities $p_{\rho}(E)$—understood here in probabilistic and operational terms rather than just as a mathematical formula—serve as quantum degrees of membership [4]. This operator-based formulation captures both crisp granules (projectors) and soft granules (non-projective effects), creating a continuum that mirrors and extends fuzzy membership functions and the boundary regions of rough sets. The effect-algebra perspective [9] gives QGC a mathematically coherent substrate grounded in quantum logic and operator theory.
@@
Although the literature contains quantum-inspired approaches in areas such as cognition, decision theory, and machine learning, no existing framework unifies operator-based granules, their algebraic properties, and computational architectures for intelligent systems. Classical GrC focuses on partitions, coverings, and approximation operators [36, 17], which cannot be straightforwardly generalized to non-commutative settings. Rough-set–based reasoning has been widely explored in intelligent systems [26, 39]; yet its underlying classical algebra remains distributive and cannot express contextuality. Recent developments emphasize granular and topological refinements [8], while effect algebras provide logical foundations without operational pipelines [9]. These gaps motivate the development of Quantum Granular Computing (QGC): a coherent framework in which granules are genuinely quantum objects, and non-commutativity is treated as an essential feature rather than a limitation.
@@
    We develop a foundational framework for effect-based granules and collect a suite of theorems characterizing their behavior, including normalization and monotonicity, Boolean islands induced by commuting families, and the evolution of granules under measurement and quantum channels via Lüders and Kraus transformations in the Heisenberg picture.
@@
    We demonstrate that optimal binary quantum decisions admit a natural interpretation as an optimal decision granule of Helstrom type, linking concepts of quantum granularity to quantum detection and estimation theory.
@@
The rest of this paper is organized as follows. Section 2 recalls the classical roots of granular computing (GrC), reviewing fuzzy, rough, interval-valued, and shadowed sets, together with mediative fuzzy logic and formal words as representative models of information granules. Section 3 introduces the quantum formalism and notation used throughout, covering finite-dimensional Hilbert spaces, density operators, projective and positive operator-valued measurements, quantum channels, and the Heisenberg picture. Section 4 defines quantum granules as effects, develops their basic algebraic and probabilistic properties, establishes Boolean-islands results for commuting families, and analyzes granular refinement under measurements and the evolution of granules under quantum channels. Section 5 formalizes Quantum Granular Decision Systems (QGDS) as an operator-valued decision architecture. Section 6 presents three reference architectures for QGC—Measurement-Driven Granular Partitioning (MDGP), Variational Effect Learning (VEL), and Hybrid Classical–Quantum pipelines (HCQ). Section 7 illustrates the framework through compact case studies on qubit granulation, two-qubit parity effects, and Helstrom-style soft decisions. Finally, Section 8 discusses implications, limitations, and future directions, and Section 9 summarizes the main contributions of the work.
@@ 2 Background
Logical and algebraic extensions of fuzzy models have been proposed to accommodate conflict and higher-order reasoning. Mediative Fuzzy Logic (MFL) provides a framework in which mediative connectives and graded entailment allow one to capture compromise and conflict resolution between granules [23, 5, 21]. More recent developments further integrate MFL with granular modeling and intelligent systems, emphasizing architectures where information granules are explicitly manipulated as computational objects [5, 32, 13]. Formal words and related constructions [18, 19] offer additional algebraic tools to describe the composition and refinement of granules in structured domains, extending the expressive power of classical GrC beyond purely numerical memberships.
@@
In this paper, we adopt a complementary perspective: instead of working with set-theoretic granules in a classical universe, we model granules as operators acting on a Hilbert space. This shift is motivated by quantum information theory, where states, measurements, and transformations are inherently operator-valued and non-commutative. Within this setting, we will reinterpret classical notions such as approximations, decision regions, and multilevel descriptions in terms of quantum effects, positive operator-valued measures (POVMs), and channels, preparing the ground for the Quantum Granular Computing framework developed in the subsequent sections.
@@ 3.1 States, measurements, and effects
Projective measurements model ideal, sharp quantum tests.
@@ 3.2 Open-system dynamics, quantum channels, and Heisenberg adjoint
that is completely positive and trace-preserving (CPTP). Complete positivity can be characterized in terms of the positivity of the so-called Choi matrix $J(\mathcal{E})$ associated with $\mathcal{E}$, defined by
@@
where $\mathrm{id}$ denotes the identity map on $\mathrm{L}(\mathcal{H}_{\mathrm{in}})$ and $|\Phi^{+}\rangle=\tfrac{1}{\sqrt{d}}\sum_{i=1}^{d}|i\rangle\otimes|i\rangle$, with $d=\dim(\mathcal{H}_{\mathrm{in}})$, is a maximally entangled state on $\mathcal{H}_{\mathrm{in}}\otimes\mathcal{H}_{\mathrm{in}}$. Choi’s theorem states that $\mathcal{E}$ is completely positive if and only if $J(\mathcal{E})\succeq 0$ [7].
@@
Physically, a channel $\mathcal{E}$ can be obtained by coupling the system to an environment, applying a global unitary, and tracing out the environment (Stinespring dilation). In the Schrödinger picture, $\mathcal{E}$ acts on states. In the Heisenberg picture, one instead considers the adjoint (dual) channel
@@
In the granular setting developed in this paper, we will use the adjoint channel $\mathcal{E}^{\dagger}$ to describe how quantum granules (effects) are transformed by noisy dynamics. This point of view will be central when analyzing the evolution of granular memberships under open-system dynamics and quantum decision processes.
@@ 3.3 Lattice of subspaces and orthomodularity
Projectors on $\mathcal{H}$ are closely related to closed subspaces: each projector $P$ is associated with its range $\operatorname{ran}(P)\subseteq\mathcal{H}$, and conversely, each closed subspace arises as the range of a unique projector. Let $\mathcal{L}(\mathcal{H})$ denote the set of all closed subspaces of $\mathcal{H}$ (equivalently, the ranges of projectors).
@@
equivalently $PQ=QP=P$. Meet and join in $\mathcal{L}(\mathcal{H})$ are defined by
@@ 3.4 Granulation, locality, and composition
which need not factorize as $E_{A}\otimes E_{B}$. Here $\mathrm{Eff}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$ denotes the set of effects on the composite system, i.e., all operators $E_{AB}$ with $0\preceq E_{AB}\preceq I_{AB}$. For entangled states $\rho_{AB}$, one may have
@@
for some joint effect $E_{AB}$, while no product interpretation in terms of separate granules on $A$ and $B$ is possible. Such global granules encode joint structure that is not reducible to independent local components and will be central to our analysis of contextuality and entanglement.
@@ 4 Quantum Granules and Foundational Results
In classical granular computing, granules are typically defined via membership functions, equivalence relations, or interval-valued descriptions. In quantum systems, these structures must be generalized to account for operator-valued information, contextuality, and incompatibility. Building on the preliminaries in Sec. 3, we introduce quantum granules as effects and establish several basic results describing their algebraic, probabilistic, and dynamical behavior. This section provides the formal core of Quantum Granular Computing (QGC).
@@ 4.1 Quantum granules
We consider effects to be the fundamental units of quantum granulation.
@@
Let $\mathcal{H}$ be a finite-dimensional Hilbert space. A quantum granule is an effect $E\in\mathrm{Eff}(\mathcal{H})$, that is, an operator satisfying
$0\preceq E\preceq I.$ (24)
Given a state $\rho\in\mathcal{D}(\mathcal{H})$, the granular membership (Born probability) of $E$ is
@@
Thus, a quantum granule is precisely an effect in the sense of quantum measurement theory [10, 4], and $p_{\rho}(E)$ is its Born-rule probability, interpreted as a granular membership degree. Projectors $P=P^{\dagger}=P^{2}$ appear as sharp granules, while general (non-projective) effects encode soft, noisy, and context-dependent acceptance regions. From the logical side, the set $\mathrm{Eff}(\mathcal{H})$ of effects, equipped with the partial sum $E\oplus F:=E+F$ whenever $E+F\preceq I$, forms an effect algebra [9], which underlies our granular constructions.

### v2

@@ Abstract
This paper introduces a formal framework for Quantum Granular Computing (QGC), which extends classical granular computing — including fuzzy, rough, and shadowed granules — to the quantum regime. Quantum granules are modeled as effects on a finite-dimensional Hilbert space, making memberships Born probabilities and embedding granulation in the standard formalism of quantum information theory. We establish structural results for effect-based granules, including normalization, monotonicity, the emergence of Boolean islands from commuting families, refinement under Lüders updates, and evolution under quantum channels via the adjoint channel in the Heisenberg picture. We connect QGC with quantum detection and estimation theory by viewing Helstrom’s minimum-error measurement for binary state discrimination as a system of Helstrom-type decision granules, i.e., soft counterparts of Bayes-optimal decision regions. Building on this, we introduce Quantum Granular Decision Systems (QGDS) and three reference architectures: Measurement-Driven Granular Partitioning, Variational Effect Learning, and Hybrid Classical–Quantum pipelines. Case studies on qubit granulation, two-qubit parity, and Helstrom-style soft decisions show that QGC reproduces fuzzy-like graded memberships and smooth decision boundaries while exposing the role of non-commutativity and entanglement. The framework provides a unified, mathematically grounded basis for operator-valued granules in quantum information processing and intelligent systems.
@@ 1 Introduction
Classical granular computing (GrC) provides a principled way to structure information through granules—coherent units such as fuzzy sets, rough sets, interval-valued descriptions, and shadowed regions—that enable reasoning under uncertainty and abstraction [43, 25]. These models play a central role in intelligent systems and have been expanded through multilevel approximation schemes and structured granulation methodologies [37, 40]. Further expressive power has been achieved through shadowed sets, interval type-2 fuzzy sets [26], and Mediative Fuzzy Logic (MFL) [23], which allow granules to encode partial compatibility, conflict, and higher-order uncertainty. Nonetheless, classical granules typically operate within Boolean or fuzzy lattices, where all granules are assumed to coexist in a single distributive structure and can always be jointly combined. This global compatibility makes it difficult to represent phenomena driven by fundamentally incompatible observations, such as non-commuting measurements in quantum theory.
@@
Quantum systems provide a natural setting in which these limitations become explicit. In quantum mechanics, states, measurements, and transformations are operators on a Hilbert space, and observable quantities generally do not commute [24, 11]. This structure gives rise to contextuality: many observables do not admit a joint description, and measurements can irreversibly disturb the system. At the same time, quantum information theory has shown that operator-valued objects such as effects, POVMs, and quantum channels offer a flexible language for soft decisions, uncertainty, and partial information [5, 11]. The lattice of closed subspaces is orthomodular and non-distributive [4], and the set of effects forms an effect algebra [10], indicating that quantum theory already contains native algebraic structures suitable for granular modeling.
@@
We therefore adopt the viewpoint that quantum granules should be modeled as effects, i.e., positive operators bounded by the identity. Effects are standard generalized events in quantum information, providing soft acceptance regions parameterized by their spectrum. Their associated Born probabilities $p_{\rho}(E)$, interpreted operationally as quantum degrees of membership, provide graded acceptance values for states [5]. This operator-based formulation captures both crisp granules (projectors) and soft granules (non-projective effects), creating a continuum that mirrors and extends fuzzy membership functions and the boundary regions of rough sets. The effect-algebra perspective [10] gives QGC a mathematically coherent substrate grounded in quantum logic and operator theory.
@@
Although the literature contains quantum-inspired approaches in areas such as cognition, decision theory, and machine learning, to the best of our knowledge no existing framework unifies operator-based granules, their algebraic properties, and computational architectures for intelligent systems. Classical GrC focuses on partitions, coverings, and approximation operators [37, 40], which cannot be straightforwardly generalized to non-commutative settings. Rough-set–based reasoning has been widely explored in intelligent systems [25, 33]; yet its underlying classical algebra remains distributive and cannot express contextuality. Recent developments emphasize granular and topological refinements [9], while effect algebras provide logical foundations without operational pipelines [10]. These gaps motivate the development of Quantum Granular Computing (QGC): a coherent framework in which granules are genuinely quantum objects, and non-commutativity is treated as an essential feature rather than a limitation.
@@
    We develop a foundational framework for effect-based granules and collect a suite of theorems characterizing their behavior, including normalization and monotonicity, Boolean islands induced by commuting families, and the evolution of granules under measurement and quantum channels via Lüders updates and the Heisenberg adjoint of Kraus channels.
@@
    We demonstrate that optimal binary quantum decisions admit a natural interpretation as an optimal decision granule of Helstrom type, linking the concepts of quantum granularity to quantum detection and estimation theory.
@@
The rest of this paper is organized as follows. Section 2 recalls the classical roots of granular computing (GrC), reviewing fuzzy, rough, interval-valued, and shadowed sets, together with mediative fuzzy logic as representative models of information granules. Section 3 introduces the quantum formalism and notation used throughout, covering finite-dimensional Hilbert spaces, density operators, projective and positive operator-valued measurements, quantum channels, and the Heisenberg picture. Section 4 defines quantum granules as effects, develops their basic algebraic and probabilistic properties, establishes results on Boolean islands for commuting families, and analyzes granular refinement under measurements and the evolution of granules under quantum channels. Section 5 formalizes Quantum Granular Decision Systems (QGDS) as an operator-valued decision architecture. Section 6 presents three reference architectures for QGC—Measurement-Driven Granular Partitioning (MDGP), Variational Effect Learning (VEL), and Hybrid Classical–Quantum pipelines (HCQ). Section 7 illustrates the framework through compact case studies on qubit granulation, two-qubit parity effects, and Helstrom-style soft decisions. Finally, Section 8 analyzes implications, limitations, and directions for further work, and Section 9 summarizes the main contributions of the paper.
@@ 2 Background
Logical and algebraic extensions of fuzzy models have been proposed to accommodate conflict and higher-order reasoning. Mediative Fuzzy Logic (MFL) provides a framework in which mediative connectives and graded entailment allow one to capture compromise and conflict resolution between granules [22, 6, 20]. More recent developments further integrate MFL with granular modeling and intelligent systems, emphasizing architectures where information granules are explicitly manipulated as computational objects [6, 32, 14]. Formal words and related constructions [17, 18] offer additional algebraic tools to describe the composition and refinement of granules in structured domains, extending the expressive power of classical GrC beyond purely numerical memberships and providing a conceptual symbolic counterpart to operator-based constructions in the quantum setting.
@@
While the present work is, to our knowledge, the first to develop a general operator-theoretic framework for Quantum Granular Computing, there are emerging contributions that combine granular ideas with quantum algorithms in more specialized settings. Yuan et al. introduce quantum granular-ball generation methods that exploit the structure of quantum computation to accelerate the construction of granular-ball summaries and propose a quantum $k$-nearest neighbors classifier (QGBkNN), showing substantial reductions in time complexity compared with classical granular-ball approaches for KNN classification [41]. Acampora and Vitiello present a hybrid granular–quantum genetic optimization scheme in which a classical granular-computing layer hierarchically reduces the continuous search space, while a quantum processor evolves the genetic population, leading to statistically significant improvements over state-of-the-art quantum evolutionary algorithms on benchmark functions under current hardware constraints [1]. These studies highlight a growing interest in leveraging granularity within quantum computing at the algorithmic level, but they do not address the foundational problem of representing information granules themselves as quantum objects, which is the focus of the present paper.
In this paper, we adopt a complementary perspective: instead of working with set-theoretic granules in a classical universe, we model granules as operators acting on a Hilbert space. This shift is motivated by quantum information theory, where states, measurements, and transformations are inherently operator-valued and non-commutative. Within this setting, we reinterpret classical notions such as approximations, decision regions, and multilevel descriptions in terms of quantum effects, positive operator-valued measures (POVMs), and channels, preparing the ground for the Quantum Granular Computing framework developed in the subsequent sections.
@@ 3.1 States, measurements, and effects
Projective measurements are often regarded as ideal, sharp quantum tests, in the sense that each outcome is represented by a projector with eigenvalues in $\{0,1\}$.
@@ 3.2 Open-system dynamics, quantum channels, and Heisenberg adjoint
that is completely positive and trace-preserving (CPTP). Complete positivity can be characterized in terms of the positivity of the Choi matrix $J(\mathcal{E})$ associated with $\mathcal{E}$, defined by
@@
where $\mathrm{id}$ denotes the identity map on $\mathrm{L}(\mathcal{H}_{\mathrm{in}})$ and $|\Phi^{+}\rangle=\tfrac{1}{\sqrt{d}}\sum_{i=1}^{d}|i\rangle\otimes|i\rangle$, with $d=\dim(\mathcal{H}_{\mathrm{in}})$, is a maximally entangled state on $\mathcal{H}_{\mathrm{in}}\otimes\mathcal{H}_{\mathrm{in}}$ for a fixed orthonormal basis $\{|i\rangle\}_{i=1}^{d}$ of $\mathcal{H}_{\mathrm{in}}$. Thus $J(\mathcal{E})$ is an operator on $\mathcal{H}_{\mathrm{out}}\otimes\mathcal{H}_{\mathrm{in}}$. Choi’s theorem states that $\mathcal{E}$ is completely positive if and only if $J(\mathcal{E})\succeq 0$ [8].
@@
Physically, a channel $\mathcal{E}$ can be obtained by coupling the system to an environment, applying a global unitary, and tracing out the environment (Stinespring dilation). In the Schrödinger picture, the dynamics is described by the evolution of states, while observables (or effects) are kept fixed: a state $\rho$ is mapped to $\mathcal{E}(\rho)$. In the Heisenberg picture, the states are kept fixed and observables or effects evolve according to the adjoint (dual) channel
@@
In the granular setting developed in this paper, we use the adjoint channel $\mathcal{E}^{\dagger}$ to describe how quantum granules (effects) are transformed by noisy dynamics. This viewpoint is central for analyzing the evolution of granular memberships under open-system dynamics and quantum decision processes.
@@ 3.3 Lattice of subspaces and orthomodularity
Projectors on $\mathcal{H}$ are closely related to closed subspaces: each projector $P$ is associated with its range $\operatorname{ran}(P)\subseteq\mathcal{H}$, and conversely, each closed subspace arises as the range of a unique orthogonal projector. Let $\mathcal{L}(\mathcal{H})$ denote the set of all closed subspaces of $\mathcal{H}$ (equivalently, the ranges of projectors).
@@
equivalently, when $PQ=QP=P$. Meet and join in $\mathcal{L}(\mathcal{H})$ are defined by
@@ 3.4 Granulation, locality, and composition
which need not factorize as $E_{A}\otimes E_{B}$. Here $\mathrm{Eff}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$ denotes the set of effects on the composite system, i.e., all operators $E_{AB}$ with $0\preceq E_{AB}\preceq I_{AB}$, where $I_{AB}=I_{A}\otimes I_{B}$. For entangled states $\rho_{AB}$, one may have
@@
for some joint effect $E_{AB}$, while no product interpretation in terms of separate granules on $A$ and $B$ is possible. Such global granules encode joint structure that is not reducible to independent local components and are central to our analysis of contextuality and entanglement.
@@ 4 Quantum Granules and Foundational Results
In classical granular computing, granules are typically defined via membership functions, equivalence relations, or interval-valued descriptions. In quantum systems, these structures must be generalized to account for operator-valued information, contextuality, and incompatibility. Building on the preliminaries in Sec. 3, we introduce quantum granules as effect-based granules and establish several basic results describing their algebraic, probabilistic, and dynamical behavior. This section provides the formal core of Quantum Granular Computing (QGC).
@@ 4.1 Quantum granules
Effects are taken as the fundamental units of quantum granulation. In analogy with classical granular models, such as fuzzy sets, rough sets, shadowed sets, MFL granules, and formal words, where a granule is represented by a universe together with one or more membership-like maps, we adopt the following definition.
@@
Let $\mathcal{H}$ be a finite-dimensional Hilbert space, let $\mathcal{D}(\mathcal{H})$ be the corresponding state space, and let $\mathrm{Eff}(\mathcal{H})$ be the set of effects on $\mathcal{H}$. A quantum granule on $\mathcal{H}$ is the triple
$\mathcal{G}=\bigl(\mathcal{D}(\mathcal{H}),E,p_{E}\bigr),$ (24)
where
  1. 1.
    $E\in\mathrm{Eff}(\mathcal{H})$ is an effect, i.e.
    $0\preceq E\preceq I,$ (25)
  2. 2.
    $p_{E}:\mathcal{D}(\mathcal{H})\to[0,1]$ is the granular membership function associated with $E$, defined by
@@
Since $\mathcal{H}$ and $\mathcal{D}(\mathcal{H})$ are fixed throughout this work, we often abbreviate the triple by writing $\mathcal{G}=(E,p_{E})$, and even identify the granule with its effect $E$, using the shorthand $p_{\rho}(E)$ for $p_{E}(\rho)$.
Mathematically, each effect $E$ thus induces a fuzzy-set–like membership function $p_{E}:\mathcal{D}(\mathcal{H})\to[0,1]$ on the quantum state space, in direct analogy with a classical fuzzy granule $(X,\mu_{A})$, where $X$ is the universe and $\mu_{A}$ assigns membership grades to its elements. Projectors $P=P^{\dagger}=P^{2}$ appear as sharp granules, with spectrum $\{0,1\}$ and $p_{P}(\rho)$ acting as a crisp acceptance indicator, while general (non-projective) effects encode soft, noisy, and context-dependent acceptance regions.
Beyond the purely fuzzy case, other classical granular formalisms can be embedded into this effect-based perspective. Rough granules, described by lower and upper approximations $(A_{L},A_{U})$, correspond in the commuting projective case to pairs of projectors $P_{L}\preceq P_{U}$, and more generally to pairs of effects $E_{L}\preceq E_{U}$ that bound the membership region for a concept. Shadowed sets and three-way decisions can be realized by finite families of effects $\{E_{\mathrm{acc}},E_{\mathrm{rej}},E_{\mathrm{und}}\}$ forming a POVM, where each effect defines an acceptance, rejection, or deferment granule at the quantum level. MFL-style mediative granules, with positive and negative components $(\mu^{+},\mu^{-})$, can be modeled by pairs of effects $(E^{+},E^{-})$ acting on the same state space, so that $p_{E^{+}}(\rho)$ and $p_{E^{-}}(\rho)$ play the role of mediative truth values; in the non-commuting case, $E^{+},E^{-}]\neq 0$ naturally encodes conflict and contextuality. Lin’s formal words, which represent structured granules built from ordered compositions of simpler units, admit a quantum counterpart as ordered tuples or products of effects (or of POVM elements), where non-commutativity records the order sensitivity of the underlying granular operations. More generally, any classical granular construct specified by a finite family of membership-like maps can be transported to the quantum setting by assigning an appropriate finite family of effects on $\mathcal{H}$ whose Born-rule memberships reproduce the intended granular semantics.
From the logical side, the set $\mathrm{Eff}(\mathcal{H})$ of effects, equipped with the partial sum $E\oplus F:=E+F$ whenever $E+F\preceq I$, forms an effect algebra [10], which provides the algebraic substrate for our granular constructions.
We record an explicit expression for the membership of pure states.
Lemma 1 (Pure-state form).
Let $\mathcal{G}=(\mathcal{D}(\mathcal{H}),E,p_{E})$ be a quantum granule on $\mathcal{H}$. For a pure state $|\rho=\ket{\psi}\!\bra{\psi}$ with $\langle\psi|\psi\rangle=1$, the granular membership is
[…]
In particular, if $E=P$ is a projector, then
[…]
Proof.
Substituting $|\rho=\ket{\psi}\!\bra{\psi}$ into (26) and using cyclicity of the trace yields
[…]
If $E=P$ is a projector, then $P^{2}=P$ and
[…]